# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Diabetic Retinopathy Severity Assessment

### v3

@@ VIII Declaration of competing interest
Layer Model SI NI II TI SO NO IO TO
$\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr
RNFL GT 0.01 0.045 0.987 0.01 0.093 0.987 0.006 0.299 0.987 0.01 0.403 0.987 0.007 0.139 0.987 0.002 0.582 0.987 0.002 0.497 0.987 0.004 0.668 0.987
Swin 0.01 0.082 0.987 0.009 0.188 0.987 0.005 0.452 0.987 0.007 0.563 0.987 0.007 0.131 0.987 0.001 0.804 0.995 0.004 0.325 0.987 -0.0 0.969 0.995
VM 0.012 0.069 0.987 0.01 0.193 0.987 0.004 0.6 0.987 0.002 0.884 0.995 0.01 0.043 0.987 0.003 0.511 0.987 0.005 0.312 0.987 0.005 0.702 0.987
GCL+IPL GT 0.0 0.987 0.996 -0.001 0.887 0.995 -0.003 0.596 0.987 -0.004 0.345 0.987 0.005 0.476 0.987 -0.002 0.77 0.987 0.005 0.489 0.987 0.004 0.52 0.987
Swin -0.001 0.876 0.995 0.0 0.962 0.995 -0.002 0.57 0.987 -0.005 0.229 0.987 0.004 0.512 0.987 0.003 0.596 0.987 0.004 0.565 0.987 0.002 0.706 0.987
VM -0.001 0.845 0.995 -0.002 0.727 0.987 -0.002 0.625 0.987 -0.005 0.277 0.987 0.002 0.739 0.987 -0.001 0.451 0.987 -0.0 0.987 0.996 0.002 0.661 0.987
INL GT -0.006 0.476 0.987 -0.018 0.116 0.987 -0.013 0.103 0.987 -0.001 0.823 0.995 0.006 0.606 0.987 -0.005 0.618 0.987 -0.008 0.313 0.987 -0.003 0.731 0.987
Swin 0.006 0.448 0.987 0.005 0.653 0.987 -0.003 0.744 0.987 0.001 0.819 0.995 0.02 0.075 0.987 0.012 0.25 0.987 0.001 0.903 0.995 0.003 0.756 0.987
VM 0.004 0.51 0.987 0.001 0.924 0.995 0.0 0.978 0.995 -0.0 0.906 0.995 0.023 0.022 0.987 0.014 0.247 0.987 0.002 0.827 0.995 0.003 0.705 0.987
OPL GT 0.008 0.139 0.987 0.003 0.59 0.987 0.002 0.46 0.987 0.003 0.251 0.987 0.03 0.028 0.987 0.007 0.46 0.987 0.003 0.635 0.987 0.004 0.575 0.987
Swin 0.014 0.039 0.987 0.003 0.701 0.987 0.004 0.673 0.987 0.009 0.079 0.987 0.016 0.176 0.987 0.007 0.618 0.987 -0.001 0.954 0.995 0.004 0.734 0.987
VM 0.005 0.23 0.987 0.003 0.511 0.987 0.004 0.494 0.987 0.004 0.18 0.987 0.024 0.034 0.987 0.021 0.207 0.987 0.003 0.755 0.987 0.003 0.76 0.987
ONL+IS GT 0.0 0.962 0.995 -0.002 0.367 0.987 0.0 0.945 0.995 -0.001 0.766 0.987 0.004 0.047 0.987 0.001 0.84 0.995 0.003 0.375 0.987 0.001 0.69 0.987
Swin 0.0 0.932 0.995 -0.002 0.584 0.987 0.001 0.781 0.992 0.001 0.676 0.987 0.005 0.06 0.987 -0.001 0.452 0.987 0.002 0.503 0.987 0.002 0.426 0.987
VM -0.001 0.747 0.987 -0.003 0.26 0.987 0.001 0.773 0.987 -0.0 0.815 0.995 0.004 0.153 0.987 -0.0 0.912 0.995 0.003 0.422 0.987 0.001 0.627 0.987
EZ GT 0.011 0.512 0.987 0.016 0.066 0.987 0.004 0.737 0.987 0.002 0.716 0.987 0.02 0.466 0.987 0.025 0.205 0.987 0.0 0.992 0.996 0.015 0.514 0.987
Swin 0.002 0.805 0.995 0.02 0.166 0.987 0.008 0.713 0.987 0.002 0.73 0.987 0.007 0.788 0.992 -0.013 0.639 0.987 -0.022 0.566 0.987 -0.002 0.949 0.995
VM 0.003 0.874 0.995 0.002 0.913 0.995 0.014 0.476 0.987 0.002 0.722 0.987 0.01 0.769 0.987 -0.033 0.288 0.987 -0.008 0.843 0.995 -0.019 0.645 0.987
OS GT 0.012 0.493 0.987 -0.023 0.211 0.987 -0.007 0.712 0.987 -0.001 0.937 0.995 0.013 0.571 0.987 -0.011 0.66 0.987 -0.01 0.678 0.987 -0.01 0.69 0.987
Swin 0.01 0.695 0.987 -0.039 0.029 0.987 -0.0 0.996 0.996 -0.001 0.97 0.995 0.021 0.508 0.987 -0.004 0.326 0.987 -0.015 0.572 0.987 -0.011 0.729 0.987
VM 0.001 0.968 0.995 -0.045 0.087 0.987 -0.018 0.468 0.987 -0.012 0.657 0.987 0.015 0.648 0.987 -0.015 0.527 0.987 -0.011 0.708 0.987 -0.024 0.476 0.987
RPE GT 0.002 0.883 0.995 -0.003 0.79 0.992 0.015 0.236 0.987 0.005 0.642 0.987 0.004 0.795 0.994 -0.005 0.691 0.987 0.03 0.049 0.987 0.024 0.082 0.987
Swin -0.0 0.967 0.995 -0.005 0.643 0.987 0.001 0.944 0.995 0.0 0.978 0.995 -0.001 0.93 0.995 -0.008 0.478 0.987 0.006 0.696 0.987 -0.008 0.669 0.987
VM 0.001 0.944 0.995 0.001 0.94 0.995 0.002 0.924 0.995 -0.006 0.718 0.987 0.013 0.5 0.987 -0.005 0.651 0.987 0.013 0.53 0.987 -0.012 0.546 0.987
Fluid GT 0.0 0.345 0.987 0.0 0.089 0.987 0.0 0.194 0.987 0.0 0.208 0.987 0.0 0.086 0.987 0.0 0.15 0.987 0.0 0.688 0.987 0.0 0.678 0.987
Swin 0.0 0.282 0.987 0.0 0.3 0.987 0.0 0.276 0.987 0.0 0.233 0.987 0.0 0.894 0.995 0.0 0.674 0.987 0.0 0.908 0.995 0.0 0.726 0.987
VM 0.0 0.217 0.987 0.0 0.135 0.987 0.0 0.151 0.987 0.0 0.191 0.987 0.0 0.001 0.168 0.0 0.483 0.987 0.0 0.538 0.987 0.0 0.468 0.987
HRF GT 0.0 0.606 0.987 0.0 0.467 0.987 -0.0 0.625 0.987 0.0 0.733 0.987 0.0 0.372 0.987 0.0 0.196 0.987 -0.0 0.725 0.987 -0.0 0.761 0.987
Swin 0.0 0.502 0.987 0.0 0.236 0.987 -0.0 0.724 0.987 0.0 0.918 0.995 0.0 0.387 0.987 0.0 0.198 0.987 -0.0 0.859 0.995 -0.0 0.725 0.987
VM 0.0 0.404 0.987 0.0 0.18 0.987 0.0 0.907 0.995 0.0 0.475 0.987 0.0 0.276 0.987 0.0 0.084 0.987 0.0 0.899 0.995 -0.0 0.97 0.995
@@
Layer Model SI NI II TI SO NO IO TO
$\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr $\bm{\beta}$ p_raw p_fdr
RNFL GT 0.004 0.596 0.996 0.009 0.329 0.996 -0.001 0.893 1.0 0.012 0.37 0.996 -0.001 0.908 1.0 -0.0 0.975 1.0 -0.001 0.9 1.0 -0.005 0.587 0.996
Swin -0.007 0.349 0.996 -0.0 0.998 1.0 -0.007 0.419 0.996 -0.015 0.383 0.996 -0.006 0.336 0.996 -0.002 0.487 0.996 -0.003 0.556 0.996 -0.015 0.077 0.996
VM -0.006 0.426 0.996 -0.003 0.774 1.0 -0.011 0.194 0.996 -0.035 0.058 0.848 -0.004 0.495 0.996 -0.004 0.511 0.996 -0.003 0.584 0.996 -0.02 0.017 0.818
GCL+IPL GT -0.004 0.439 0.996 -0.003 0.575 0.996 -0.006 0.222 0.996 -0.002 0.724 1.0 -0.012 0.089 0.996 -0.012 0.106 0.996 -0.007 0.273 0.996 -0.006 0.398 0.996
Swin -0.0 0.975 1.0 0.0 0.951 1.0 -0.005 0.277 0.996 -0.0 0.949 1.0 -0.008 0.252 0.996 -0.004 0.582 0.996 -0.005 0.474 0.996 -0.0 0.981 1.0
VM -0.0 0.917 1.0 -0.001 0.861 1.0 -0.004 0.439 0.996 0.0 0.975 1.0 -0.006 0.36 0.996 -0.007 0.357 0.996 -0.006 0.391 0.996 -0.002 0.807 1.0
INL GT -0.008 0.392 0.996 -0.006 0.581 0.996 -0.022 0.049 0.84 -0.011 0.326 0.996 -0.008 0.44 0.996 -0.002 0.88 1.0 -0.014 0.242 0.996 -0.016 0.178 0.996
Swin -0.005 0.623 0.996 -0.009 0.457 0.996 -0.01 0.357 0.996 -0.002 0.827 1.0 -0.005 0.659 1.0 -0.004 0.792 1.0 -0.014 0.247 0.996 -0.006 0.595 0.996
VM 0.003 0.813 1.0 -0.003 0.749 1.0 -0.011 0.271 0.996 0.005 0.62 0.996 -0.004 0.672 1.0 0.002 0.901 1.0 -0.013 0.253 0.996 -0.005 0.64 0.996
OPL GT 0.005 0.371 0.996 0.004 0.408 0.996 0.004 0.295 0.996 0.004 0.358 0.996 0.004 0.756 1.0 0.002 0.91 1.0 -0.006 0.619 0.996 -0.0 0.967 1.0
Swin 0.004 0.562 0.996 0.007 0.291 0.996 0.004 0.469 0.996 0.002 0.694 1.0 0.0 0.997 1.0 -0.007 0.721 1.0 -0.01 0.387 0.996 -0.006 0.519 0.996
VM 0.006 0.238 0.996 0.008 0.06 0.848 0.004 0.326 0.996 0.003 0.325 0.996 0.003 0.744 1.0 0.014 0.403 0.996 -0.004 0.738 1.0 0.002 0.821 1.0
ONL+IS GT -0.001 0.701 1.0 -0.003 0.443 0.996 -0.004 0.15 0.996 -0.003 0.299 0.996 -0.001 0.639 0.996 0.0 0.89 1.0 -0.002 0.612 0.996 -0.001 0.626 0.996
Swin 0.002 0.496 0.996 -0.002 0.559 0.996 -0.004 0.284 0.996 0.001 0.822 1.0 -0.0 0.874 1.0 0.002 0.72 1.0 -0.002 0.647 0.996 -0.0 0.583 0.996
VM 0.002 0.595 0.996 -0.001 0.82 1.0 -0.002 0.482 0.996 -0.001 0.85 1.0 -0.0 0.942 1.0 0.001 0.773 1.0 -0.002 0.526 0.996 -0.001 0.733 1.0
EZ GT 0.024 0.247 0.996 0.005 0.467 0.996 0.01 0.418 0.996 0.002 0.675 1.0 0.004 0.899 1.0 0.06 0.028 0.818 0.023 0.498 0.996 0.034 0.124 0.996
Swin 0.024 0.426 0.996 0.018 0.494 0.996 0.002 0.856 1.0 -0.002 0.861 1.0 0.008 0.832 1.0 0.064 0.167 0.996 0.024 0.617 0.996 -0.011 0.216 0.996
VM 0.012 0.681 1.0 0.003 0.861 1.0 -0.006 0.644 0.996 -0.004 0.737 1.0 0.008 0.788 1.0 0.022 0.512 0.996 0.015 0.688 1.0 0.019 0.412 0.996
OS GT 0.041 0.059 0.848 0.039 0.101 0.996 -0.011 0.647 0.996 0.027 0.262 0.996 0.054 0.048 0.84 0.054 0.09 0.996 0.059 0.044 0.84 0.066 0.031 0.818
Swin 0.014 0.617 0.996 0.033 0.321 0.996 0.024 0.492 0.996 0.023 0.44 0.996 0.074 0.021 0.818 0.071 0.049 0.84 0.074 0.033 0.818 0.025 0.325 0.996
VM 0.038 0.128 0.996 0.032 0.211 0.996 0.036 0.211 0.996 0.039 0.134 0.996 0.087 0.003 0.352 0.09 0.008 0.504 0.107 0.001 0.336 0.001 0.951 1.0
RPE GT -0.014 0.432 0.996 -0.014 0.335 0.996 -0.018 0.266 0.996 -0.0 0.996 1.0 0.012 0.53 0.996 0.002 0.902 1.0 0.013 0.497 0.996 0.011 0.601 0.996
Swin 0.006 0.691 1.0 -0.02 0.174 0.996 -0.037 0.034 0.818 -0.005 0.749 1.0 -0.002 0.937 1.0 -0.003 0.897 1.0 -0.0 0.99 1.0 0.014 0.547 0.996
VM 0.0 0.638 0.996 -0.026 0.189 0.996 -0.055 0.004 0.352 0.0 0.639 0.996 -0.006 0.778 1.0 0.003 0.434 0.996 -0.032 0.198 0.996 0.0 0.638 0.996
Fluid GT 0.0 0.383 0.996 0.0 0.55 0.996 0.0 0.112 0.996 0.0 0.421 0.996 0.0 0.284 0.996 0.0 0.36 0.996 0.0 0.311 0.996 0.0 0.646 0.996
Swin 0.0 0.645 0.996 0.0 0.424 0.996 0.0 0.165 0.996 0.0 0.531 0.996 0.0 0.62 0.996 0.0 0.423 0.996 -0.0 0.784 1.0 -0.0 0.536 0.996
VM 0.0 0.535 0.996 0.0 0.589 0.996 0.0 0.323 0.996 0.0 0.567 0.996 0.0 0.42 0.996 0.0 0.453 0.996 0.0 0.654 1.0 0.0 0.953 1.0
HRF GT 0.0 0.739 1.0 -0.0 0.705 1.0 0.0 0.784 1.0 -0.0 0.874 1.0 -0.0 0.964 1.0 -0.0 0.997 1.0 -0.0 0.923 1.0 -0.0 0.236 0.996
Swin 0.0 0.82 1.0 0.0 0.993 1.0 0.0 0.779 1.0 -0.0 0.392 0.996 0.0 0.885 1.0 -0.0 0.942 1.0 -0.0 0.76 1.0 -0.0 0.115 0.996
VM 0.0 0.865 1.0 -0.0 1.0 1.0 0.0 0.922 1.0 -0.0 0.703 1.0 -0.0 0.977 1.0 0.0 0.762 1.0 -0.0 0.698 1.0 -0.0 0.208 0.996

### v4

@@ C SI NI II TI SO NO IO TO
C SI NI II TI SO NO IO TO
$\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR
RNFL GT / / / 0.019 0.000 0.002 0.017 0.015 0.030 0.014 0.013 0.027 0.038 0.000 0.003 0.013 0.004 0.014 0.011 0.015 0.030 0.007 0.103 0.110 0.027 0.000 0.003
Swin / / / 0.011 0.052 0.072 0.016 0.050 0.070 0.014 0.020 0.036 0.037 0.001 0.005 0.008 0.095 0.106 0.002 0.560 0.381 0.007 0.093 0.106 0.030 0.000 0.002
VM / / / 0.019 0.006 0.016 0.024 0.009 0.020 0.014 0.036 0.055 0.034 0.016 0.032 0.008 0.131 0.128 0.009 0.119 0.119 0.004 0.369 0.282 0.023 0.024 0.041
GCL+IPL GT / / / 0.004 0.363 0.280 0.004 0.367 0.282 0.003 0.582 0.387 0.006 0.216 0.189 0.003 0.680 0.437 0.012 0.092 0.106 0.012 0.102 0.110 0.003 0.605 0.399
Swin / / / 0.002 0.628 0.409 0.002 0.730 0.463 -0.000 0.943 0.570 0.003 0.501 0.353 0.001 0.849 0.516 0.004 0.582 0.387 0.011 0.103 0.110 -0.000 0.975 0.588
VM / / / 0.002 0.641 0.416 0.005 0.359 0.278 0.001 0.800 0.494 0.004 0.406 0.302 0.008 0.173 0.157 0.011 0.124 0.123 0.019 0.005 0.016 0.005 0.374 0.285
INL GT / / / 0.010 0.248 0.213 0.020 0.100 0.110 0.010 0.470 0.338 0.005 0.564 0.382 0.025 0.005 0.016 0.035 0.000 0.001 0.029 0.000 0.001 0.015 0.058 0.077
Swin / / / 0.013 0.113 0.115 0.037 0.002 0.008 0.030 0.005 0.015 0.004 0.506 0.354 0.022 0.016 0.032 0.039 0.004 0.014 0.037 0.000 0.002 0.012 0.156 0.144
VM / / / 0.012 0.049 0.070 0.035 0.000 0.001 0.035 0.000 0.002 0.011 0.086 0.103 0.027 0.001 0.003 0.048 0.000 0.000 0.039 0.000 0.000 0.020 0.019 0.036
OPL GT / / / 0.001 0.751 0.475 0.004 0.350 0.276 0.004 0.152 0.143 0.001 0.817 0.501 0.030 0.006 0.016 0.017 0.072 0.089 0.014 0.019 0.036 0.015 0.060 0.078
Swin / / / 0.015 0.033 0.052 0.033 0.000 0.000 0.030 0.000 0.000 0.009 0.096 0.106 0.028 0.004 0.014 0.069 0.000 0.000 0.037 0.000 0.001 0.015 0.068 0.085
VM / / / 0.002 0.516 0.359 0.015 0.002 0.010 0.011 0.008 0.020 0.001 0.720 0.459 0.026 0.001 0.006 0.069 0.000 0.000 0.028 0.000 0.002 0.014 0.054 0.072
ONL+IS GT / / / 0.001 0.624 0.408 -0.003 0.271 0.226 -0.004 0.091 0.106 -0.006 0.004 0.014 0.004 0.011 0.022 0.005 0.088 0.104 0.004 0.205 0.182 0.001 0.818 0.501
Swin / / / 0.001 0.556 0.381 -0.005 0.125 0.123 -0.004 0.115 0.116 -0.006 0.037 0.056 0.006 0.005 0.015 0.005 0.129 0.127 0.004 0.132 0.128 0.001 0.560 0.381
VM / / / 0.002 0.337 0.270 -0.003 0.291 0.239 -0.003 0.203 0.181 -0.004 0.139 0.133 0.006 0.003 0.014 0.005 0.111 0.115 0.005 0.064 0.083 0.002 0.355 0.277
EZ GT 0.003 0.089 0.104 0.027 0.073 0.089 0.019 0.004 0.015 0.018 0.073 0.089 0.007 0.105 0.110 0.066 0.006 0.017 0.062 0.000 0.003 0.060 0.020 0.036 0.062 0.009 0.020
Swin 0.004 0.165 0.151 0.010 0.264 0.222 0.034 0.009 0.020 0.016 0.105 0.110 0.007 0.181 0.163 0.060 0.010 0.022 0.076 0.009 0.020 0.089 0.011 0.022 0.059 0.056 0.075
VM 0.005 0.038 0.057 0.033 0.082 0.099 0.032 0.004 0.014 0.021 0.038 0.057 0.008 0.110 0.115 0.101 0.000 0.002 0.094 0.000 0.002 0.100 0.003 0.014 0.044 0.039 0.059
OS GT 0.000 0.990 0.594 -0.018 0.249 0.213 -0.019 0.225 0.195 -0.010 0.556 0.381 -0.025 0.092 0.106 -0.011 0.590 0.391 -0.023 0.317 0.258 -0.008 0.709 0.454 -0.006 0.782 0.491
Swin -0.036 0.006 0.016 -0.044 0.031 0.051 -0.048 0.028 0.047 -0.041 0.076 0.092 -0.043 0.032 0.051 -0.015 0.617 0.405 -0.041 0.151 0.143 -0.035 0.224 0.195 -0.025 0.261 0.220
VM -0.031 0.007 0.017 -0.015 0.476 0.340 -0.019 0.353 0.277 -0.019 0.388 0.291 -0.031 0.133 0.128 0.034 0.252 0.214 0.017 0.580 0.387 0.009 0.763 0.480 0.014 0.649 0.419
RPE GT -0.023 0.007 0.017 0.003 0.796 0.494 -0.021 0.050 0.070 -0.009 0.496 0.351 -0.012 0.319 0.258 0.022 0.151 0.143 -0.013 0.315 0.258 0.015 0.344 0.273 0.012 0.408 0.302
Swin -0.023 0.008 0.020 -0.016 0.157 0.144 -0.020 0.103 0.110 -0.007 0.568 0.383 -0.028 0.059 0.078 0.012 0.412 0.302 -0.013 0.413 0.302 0.013 0.459 0.333 0.012 0.554 0.381
VM 0.000 0.006 0.017 0.001 0.006 0.016 -0.003 0.849 0.516 0.004 0.797 0.494 0.000 0.006 0.016 0.031 0.067 0.085 -0.004 0.338 0.270 0.015 0.469 0.338 0.000 0.006 0.016
Fluid GT 0.000 0.020 0.036 0.000 0.155 0.144 0.000 0.010 0.021 0.000 0.001 0.003 0.000 0.115 0.116 0.000 0.023 0.040 0.000 0.043 0.062 0.000 0.024 0.041 0.000 0.066 0.084
Swin 0.000 0.044 0.064 0.000 0.388 0.291 0.000 0.004 0.015 0.000 0.000 0.002 0.000 0.272 0.226 0.000 0.112 0.115 0.000 0.027 0.044 0.000 0.180 0.163 0.000 0.282 0.234
VM 0.000 0.042 0.062 0.000 0.494 0.351 0.000 0.026 0.044 0.000 0.004 0.014 0.000 0.412 0.302 0.000 0.032 0.052 0.000 0.336 0.270 0.000 0.216 0.189 0.000 0.507 0.354
HRF GT 0.000 0.001 0.004 0.000 0.004 0.015 0.000 0.053 0.072 0.000 0.001 0.003 0.000 0.041 0.060 0.000 0.093 0.106 0.000 0.005 0.015 0.000 0.024 0.041 0.000 0.052 0.072
Swin 0.000 0.000 0.002 0.000 0.018 0.035 0.000 0.001 0.006 0.000 0.000 0.001 0.000 0.000 0.002 0.000 0.381 0.289 0.000 0.000 0.003 0.000 0.004 0.014 0.000 0.095 0.106
VM 0.000 0.102 0.110 0.000 0.426 0.310 0.000 0.005 0.015 0.000 0.022 0.040 0.000 0.009 0.020 0.000 0.797 0.494 0.000 0.007 0.017 0.000 0.053 0.072 0.000 0.215 0.189
@@
C SI NI II TI SO NO IO TO
$\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR $\beta$ praw pFDR
RNFL GT / / / -0.004 0.526 0.962 -0.004 0.584 0.991 -0.006 0.436 0.929 0.001 0.922 0.993 -0.004 0.590 0.991 -0.008 0.104 0.647 -0.005 0.297 0.838 0.022 0.015 0.478
Swin / / / -0.004 0.653 0.993 -0.001 0.872 0.993 -0.006 0.500 0.959 0.005 0.769 0.993 -0.003 0.673 0.993 -0.002 0.659 0.993 -0.003 0.636 0.993 0.025 0.004 0.478
VM / / / -0.008 0.345 0.880 -0.008 0.406 0.929 -0.011 0.164 0.688 0.004 0.841 0.993 -0.005 0.488 0.959 -0.006 0.216 0.734 -0.006 0.299 0.838 0.027 0.004 0.478
GCL+IPL GT / / / -0.013 0.015 0.478 -0.007 0.209 0.723 -0.007 0.242 0.762 -0.008 0.140 0.688 -0.009 0.336 0.873 -0.008 0.343 0.880 -0.016 0.057 0.616 -0.011 0.099 0.647
Swin / / / -0.012 0.027 0.616 -0.003 0.524 0.962 -0.007 0.249 0.764 -0.008 0.129 0.688 -0.006 0.571 0.991 0.000 0.961 0.994 -0.010 0.205 0.723 -0.013 0.063 0.616
VM / / / -0.010 0.050 0.616 -0.007 0.209 0.723 -0.005 0.400 0.929 -0.006 0.221 0.743 -0.008 0.366 0.918 -0.001 0.506 0.959 -0.014 0.057 0.616 -0.010 0.134 0.688
INL GT / / / 0.011 0.242 0.762 0.013 0.173 0.688 0.007 0.432 0.929 0.007 0.078 0.647 0.006 0.751 0.993 0.023 0.198 0.723 -0.009 0.586 0.991 0.004 0.738 0.993
Swin / / / 0.001 0.928 0.993 0.016 0.177 0.693 0.008 0.464 0.955 0.008 0.050 0.616 -0.001 0.922 0.993 0.017 0.169 0.688 -0.004 0.823 0.993 0.003 0.826 0.993
VM / / / 0.000 0.973 0.994 0.013 0.179 0.693 0.010 0.229 0.757 0.006 0.163 0.688 -0.002 0.902 0.993 0.021 0.237 0.762 0.003 0.851 0.993 -0.001 0.923 0.993
OPL GT / / / -0.005 0.431 0.929 0.000 0.973 0.994 0.000 0.988 0.994 -0.003 0.476 0.959 -0.001 0.948 0.993 -0.002 0.932 0.993 0.023 0.160 0.688 -0.002 0.857 0.993
Swin / / / 0.001 0.942 0.993 0.004 0.548 0.971 0.005 0.422 0.929 -0.002 0.713 0.993 -0.002 0.899 0.993 0.015 0.312 0.838 0.003 0.861 0.993 0.007 0.599 0.991
VM / / / 0.002 0.660 0.993 0.002 0.670 0.993 0.003 0.435 0.929 -0.000 0.866 0.993 0.003 0.843 0.993 0.020 0.303 0.838 0.012 0.543 0.971 0.003 0.768 0.993
ONL+IS GT / / / -0.002 0.400 0.929 0.001 0.873 0.993 -0.006 0.150 0.688 -0.006 0.036 0.616 -0.001 0.736 0.993 0.004 0.249 0.764 -0.001 0.887 0.993 -0.002 0.500 0.959
Swin / / / 0.000 0.954 0.993 0.003 0.448 0.929 -0.002 0.593 0.991 -0.006 0.101 0.647 0.001 0.850 0.993 0.000 0.900 0.993 -0.001 0.890 0.993 0.000 0.059 0.616
VM / / / -0.001 0.674 0.993 0.001 0.715 0.993 -0.002 0.595 0.991 -0.005 0.100 0.647 0.000 0.928 0.993 0.003 0.273 0.809 0.001 0.810 0.993 -0.000 0.949 0.993
EZ GT 0.001 0.757 0.993 -0.014 0.547 0.971 0.003 0.705 0.993 0.002 0.891 0.993 0.003 0.606 0.991 -0.011 0.770 0.993 0.006 0.835 0.993 -0.021 0.528 0.962 -0.009 0.689 0.993
Swin -0.000 0.950 0.993 -0.003 0.912 0.993 -0.001 0.975 0.994 -0.001 0.965 0.994 0.011 0.693 0.993 -0.043 0.312 0.838 -0.026 0.515 0.959 -0.079 0.100 0.647 -0.002 0.804 0.993
VM -0.001 0.886 0.993 -0.004 0.903 0.993 0.007 0.780 0.993 -0.006 0.875 0.993 0.002 0.912 0.993 -0.006 0.870 0.993 -0.013 0.725 0.993 -0.013 0.772 0.993 -0.033 0.367 0.918
OS GT -0.027 0.115 0.684 -0.052 0.099 0.647 -0.059 0.081 0.647 -0.052 0.121 0.688 -0.053 0.065 0.616 -0.074 0.107 0.650 -0.046 0.312 0.838 -0.060 0.134 0.688 -0.078 0.063 0.616
Swin -0.015 0.593 0.991 -0.017 0.641 0.993 -0.028 0.210 0.723 -0.012 0.745 0.993 -0.023 0.554 0.974 -0.055 0.206 0.723 -0.004 0.442 0.929 -0.060 0.098 0.647 -0.062 0.157 0.688
VM -0.015 0.626 0.993 -0.030 0.405 0.929 -0.036 0.276 0.809 -0.025 0.477 0.959 -0.029 0.443 0.929 -0.061 0.138 0.688 -0.015 0.606 0.991 -0.043 0.320 0.849 0.016 0.144 0.688
RPE GT -0.012 0.303 0.838 -0.028 0.084 0.647 -0.020 0.173 0.688 -0.023 0.140 0.688 -0.020 0.165 0.688 -0.015 0.409 0.929 -0.012 0.488 0.959 -0.026 0.142 0.688 -0.016 0.421 0.929
Swin -0.012 0.405 0.929 0.003 0.864 0.993 0.000 0.989 0.994 0.017 0.406 0.929 -0.002 0.932 0.993 -0.002 0.930 0.993 0.003 0.828 0.993 -0.013 0.502 0.959 -0.006 0.783 0.993
VM -0.008 0.651 0.993 -0.015 0.509 0.959 -0.001 0.937 0.993 0.009 0.671 0.993 -0.008 0.744 0.993 -0.000 0.992 0.994 -0.006 0.705 0.993 0.003 0.922 0.993 -0.025 0.335 0.873
Fluid GT 0.000 0.013 0.478 0.000 0.006 0.478 0.000 0.031 0.616 0.000 0.023 0.616 0.000 0.039 0.616 0.000 0.515 0.959 0.000 0.191 0.723 0.000 0.170 0.688 0.000 0.170 0.688
Swin 0.000 0.013 0.478 0.000 0.088 0.647 0.000 0.059 0.616 0.000 0.104 0.647 0.000 0.064 0.616 -0.000 0.203 0.723 0.000 0.240 0.762 0.000 0.675 0.993 -0.000 0.951 0.993
VM 0.000 0.012 0.478 0.000 0.031 0.616 0.000 0.042 0.616 0.000 0.031 0.616 0.000 0.049 0.616 0.000 0.910 0.993 0.000 0.267 0.800 0.000 0.505 0.959 0.000 0.994 0.994
HRF GT 0.000 0.697 0.993 0.000 0.447 0.929 0.000 0.153 0.688 -0.000 0.633 0.993 -0.000 0.731 0.993 -0.000 0.261 0.792 -0.000 0.747 0.993 -0.000 0.987 0.994 -0.000 0.847 0.993
Swin 0.000 0.566 0.988 0.000 0.688 0.993 0.000 0.080 0.647 0.000 0.489 0.959 0.000 0.427 0.929 -0.000 0.882 0.993 0.000 0.390 0.929 0.000 0.418 0.929 0.000 0.616 0.993
VM 0.000 0.615 0.993 0.000 0.762 0.993 0.000 0.082 0.647 0.000 0.397 0.929 0.000 0.301 0.838 -0.000 0.821 0.993 0.000 0.545 0.971 0.000 0.719 0.993 0.000 0.641 0.993